3. A wastewater treatment plant adds $700,000 to the cost of a computer chip manufacturing facility interest rate is 4%, and the end-of-life scrap value is $40,000? (6 points) due to capital expenditure for wastewater treatment? (3 points) a) What is the annualized capital cost (U) of this plant if its lifetime is 15 years, the b) If the facility manufactures 1 million chips per year, what is the added cost per chip
Answer #1

Answer:

Given

Plant Cost P=$700,000

Life cycle n=15 years

Salvage value S=$40,000

Interest Rate r=4%

A)

Annualized Capital cost U=Equivalent annual cost of initial investment P for 15 years-Equivalent annual cost of Salvage value S for 15 years

Equivalent annual cost of initial investment P for 15 years=P*r/(1-(1+r)^-n)=$62958.77

Equivalent annual cost of Salvage Value S for 15 years=S*r/((1+r)^n-1)=$3597.64

U=62958.77-3597.64=$59361.12

B)

Production C=1,000,000 chips per year

Added cost per chip= U/C=59361.12/1000000=$0.059 per chip